(0.4 * 10^-6)(0.7 * 10^2)
Could you please solve and show the work?

Answer by richwmiller(17219)

We use * for multiplication on computers.
2800 is the result
10^6/10^(-2)=10^4
.4*.7=.28=2.8*10^-1
so now we have
2.8*10^-1*10^4=
2.8*10^3=
2800